Be sure to answer all parts. An aqueous solution of a strong base has pH 10.88 at 25°C. Calculate the concentration of base in the solution: (a) if the base is LiOH. [LiOH] = M (b) if the base is Ba(OH)2. [Ba(OH)2] = × 10 M (Enter your answer in scientific notation.)
Answer #1


a) The pH of LiOH solution =10.88 pOH = 14.00 -PH = 14.00 - 10.88 = 3.12 POH = -log[OH-] [OH-] = 10 DOH = 10-312 = 0.00076 M